Instructions for ACL 2005 Workshop on Software Proceedings

Martin Jansche Center for Computational Learning Systems Columbia University New York, NY 10027, USA

Abstract

This document contains special instructions for preparing papers for inclusion in the proceedings of the ACL 2005 Workshop on Software.

1 Obligatory

Important: Follow the formatting guidelines for the Main Session, unless otherwise noted.

However, the workshop proceedings differ in several important aspects. First, there are no page limitations. However, I would strongly advise you to keep your papers under 20 pages total. Please format papers according to standard procedure: floating figures and tables should occur close to where they are referenced (not in an appendix). If you want to include source code for software that you’ve written, please submit additional files – don’t include complete source code in the paper’s appendices.

A second difference is page size. Papers should be formatted for a page size of 595 by 792 PostScript points (“big points” in TEX terminology). This is the size of the largest rectangle that will fit in both US Letter paper and ISO A4 paper. (This does not seem to work if you work in a traditional DVI → PostScript → PDF chain; in this case, please use US Letter paper).

Third, please use the mathptmx package instead of the obsolete times package. See the source code of this file for an example.

Fourth, feel free to use color in figures. To include figures, please use \includegraphics from the standard L^{A}TEX package graphicx.

systolic (mm Hg) income (kUSD)
120  80
200  160

Table 1: A floating table

You will need to submit high-quality PDF files. This almost invariable means using a recent version of pdfL^{A}TEX to prepare PDF files. I recommend the teTEX distribution version 3.0 or higher. Make sure that all fonts (except for possibly the PDF “Base 14” fonts) are embedded as Type 1 or TrueType fonts in the PDF file. You can check which fonts are included using Acrobat or the pdffonts utility that comes with xpdf.

2 Optional

You may consider using the hyperref package if you want to use internal or external links (this is completely optional). See the source code of this file for an example. It allows you to use links to sections and tables, like this one: Table 1.

If you do use hyperref you cannot use ACL’s acl.bst BibTEX style. Instead, use the natbib package together with the aclnat.bst style file provided by me. This allows you to create automatically hyperlinked references like this: “It has been claimed (?)… ” or “? claims… ”

References

Martin Kay. 1979. Syntactic processes. In Proceedings of the 17th Annual Meeting of the Association for Computational Linguistics, pages 1–2. ACL Anthology P79-1001.